# Open Panoramic Segmentation

### v1

@@ Abstract
Panoramic images, capturing a 360° field of view (FoV), encompass omnidirectional spatial information crucial for scene understanding. However, it is not only costly to obtain training-sufficient dense-annotated panoramas but also application-restricted when training models in a close-vocabulary setting. To tackle this problem, in this work, we define a new task termed Open Panoramic Segmentation (OPS), where models are trained with FoV-restricted pinhole images in the source domain in an open-vocabulary setting while evaluated with FoV-open panoramic images in the target domain, enabling the zero-shot open panoramic semantic segmentation ability of models. Moreover, we propose a model named OOOPS with a Deformable Adapter Network (DAN), which significantly improves zero-shot panoramic semantic segmentation performance. To further enhance the distortion-aware modeling ability from the pinhole source domain, we propose a novel data augmentation method called Random Equirectangular Projection (RERP) which is specifically designed to address object deformations in advance. Surpassing other state-of-the-art open-vocabulary semantic segmentation approaches, a remarkable performance boost on three panoramic datasets, WildPASS, Stanford2D3D, and Matterport3D, proves the effectiveness of our proposed OOOPS model with RERP on the OPS task, especially +2.2% on outdoor WildPASS and +2.4% mIoU on indoor Stanford2D3D. The code will be available at OPS.
@@ 2.1 Open-vocabulary Semantic Segmentation
Motivated by the success of vision-language pre-training models, Open-Vocabulary Semantic Segmentation (OVSS) has been recently explored to achieve open-set pixel-wise understanding [32, 73, 19, 12, 56, 74, 61, 45, 35, 40, 20, 41, 64, 33, 7, 11, 59, 35, 62, 67, 30, 38]. Most works leverage the open-vocabulary capabilities of CLIP [51] to reason about relations between dense pixel semantics and arbitrary class labels [9, 33, 64, 39, 74, 12]. In [91, 81], multiple datasets are unified by transferring heterogeneous labels into text embeddings for training to achieve OVSS. In [36, 70], image-caption datasets are employed for learning alignment between visual regions and text entities. To reduce the computational costs of using a frozen pre-trained CLIP model for identifying novel classes, latest studies [72, 22] have also proposed efficient architectures for fast OVSS. SAN [72, 71] has designed an asymmetric input resolution strategy to decouple mask proposal prediction and attention bias prediction applied to CLIP for mask class recognition. Dao et al. [11] propose a class enhancement loss for OVSS by using pseudo labels. Pandy et al. [49] propose a language-guided benchmark for weakly supervised OVSS. Global knowledge distillation [22] is introduced to impose constraints between the visual queries and the corresponding region embeddings produced via CLIP. Furthermore, open-vocabulary panoptic segmentation has been tackled in [12, 69, 8, 13, 83, 46, 74, 69, 35] to render universal image understanding. Garg et al. [18] propose open-world visual navigation based on panoramic semantic segmentation. Diffusion-based methods are leveraged by [59] for panoptic OVSS. In this work, we focus on the open-vocabulary zero-shot learning ability on panoramas instead of pinhole images discussed in the aforementioned works.
@@ 2.3 Deformable Neural Networks
In the realm of deformable neural networks, notable advancements have been made through various iterations. The pioneering work [10] introduces the concept of deformable convolution, laying the foundation for subsequent developments. Building upon this foundation, DCNv2 [92] further refines deformable convolution by introducing a modulation mechanism that expands the scope of deformation modeling. The evolution of deformable neural networks continues with the introduction of DCNv3 [60], which combines the group convolution with DCNv2 for the sake of imitating multi-head mechanism of Transformer [58]. DCNv4 [68] aims to speedup DCNv3, achieving a more efficient model. Additionally, Trans4PASS [85, 86] proposes Deformable Patch Embedding (DPE) and Deformable MLP (DMLP) to handle the distortion that occurs in panoramas. Incorporating the principles of deformable convolution into vision transformers [14], Xia et al. propose DeformableViT [65, 66], extending the applicability of deformable neural networks to the domain of transformer architectures. Shi et al. [55] propose deformable attention to achieve better results in open-vocabulary dense prediction. After rethinking a series of DCN designs, we propose DAO to improve the model’s ability to handle image distortion in the OPS task.
@@ 3.1 Open Panoramic Segmentation
Figure 2: Overview of the OOOPS model architecture. It consists of a frozen CLIP model and a Deformable Adapter Network (DAN) which includes Transformer Layers and the proposed DAO. Feature fusion occurs between the intermediate layers of CLIP and DAN. One of two outputs from DAN is the mask proposals, and the other one is used as deep supervised guidance for CLIP to produce proposal logits. During the training pinhole images are fed into OOOPS producing mask proposals and proposal logits for the loss calculation. During the inference panoramas are fed into OOOPS to produce the segmentation predictions.
@@ 3.3 Deformable Adapter Network
The Deformable Adapter Network (DAN) is a combination of multiple Transformer [58] Layers and the proposed DAO. Since distortion exists in panoramic images, which is a big challenge when utilizing informative panoramas, we delve deeper into the deformable convolution design [60] and the sampling method like APES [63] and propose DAO to tackle the image distortions and object deformations of panoramas. To provide a detailed rethinking process of the deformable convolution design, we introduce the preliminary of the DCN series as follows.
@@
where $\mathbf{m}_{k}$ is the learnable modulation scalar for the $k$-th location. Inspired by the Multi-Head Self-Attention (MHSA) mechanism of Transformer [58], DCNv3 [60] is proposed with a grouping operation for the sake of imitating MHSA, further boosting the deformation-aware ability of DCNv2. DCNv3 can be addressed with the following formula:
@@
Deformable Adapter Operator (DAO). It has been shown that DCNv3 and DCNv4 effectively address deformations in pinhole images. However, when it comes to distortions in panoramas, they fall short of meeting the distortion-aware requirements. This is illustrated by the visualization of Fig. 6 in Sec. 4.5. Therefore, DAO is proposed to tackle the distortion problem in panoramic images using the following expression:
@@ 3.4 Random Equirectangular Projection
The Equirectangular Projection [52] (ERP) maps meridians to vertical straight lines of constant spacing and circles of latitude to horizontal straight lines of constant spacing. ERP is one of the most common methods for mapping a globe into a panoramic plane, which transforms spherical coordinates into planar coordinates expressed as follows:
@@ 4.1 Datasets
In this work, we include four datasets ranging from the pinhole to the panoramic domains and from the indoor to the outdoor scenes. We train our model using the pinhole general-object COCOStuff-164k [5] dataset while evaluating the model performance on three different panoramic datasets: panoramic street-view WildPASS dataset [79], panoramic indoor-scene Stanford2D3D dataset [2] and panoramic indoor-scene Matterport3D dataset [6].
@@
Matterport3D. The panoramic indoor dataset includes $10,800$ panoramic views from $194,400$ RGB-D images. Following the data preprocessing pipeline of 360BEV [57], a subset of $772$ images in the size of $512{\times}1024$ with $20$ classes is used to evaluate the segmentation performance.
@@ 4.4 Comparison with the State of the Art
Table 1: Results on the WildPASS dataset. Methods of both close- and open-vocabulary panoramic semantic segmentation are compared. The mIoU, open mIoU, and per-class IoU are in percentage (%). #PARAMs is the number of learnable parameters in millions.
@@
Table 2: Results on the Stanford2D3D dataset. Methods of both close- and open-vocabulary panoramic semantic segmentation are compared. The mIoU, open mIoU, and per-class IoU are in percentage (%). #PARAMs is the number of learnable parameters in millions.
@@
Table 3: Results on the Matterport3D dataset. Methods of both close- and open-vocabulary panoramic semantic segmentation are compared. The mIoU, open mIoU, and per-class IoU are in percentage (%). #PARAMs is the number of learnable parameters in millions.
@@ 4.5 Ablation Study
We have conducted an ablation study to showcase the effectiveness of both DAO and RERP. In particular, we replace DAO in OOOPS, as illustrated in Fig. 2, with DCN [10], DCNv2 [92], DCNv3 [60] and DCNv4 [68]. When referring to the component substitution without data augmentation in Table 5, it is clear that our DAO surpasses other counterparts with a maximum performance gain of +1.4% mIoU. As for the component substitution with ERP, all models achieve a boost in the range of ${+}0.3\%$ to ${+}0.7\%$ compared with the one without data augmentation. The gain is even bigger considering the component substitution with RERP, achieving ${+}1.0\%$ mIoU. It is worth noting that RERP is able to boost the model performance stably among all models, e.g., ${+}1.0\%$ performance enhancement in DCNv3 substitution. The best model performance is achieved by the combination of DAO and RERP, yielding a mIoU of 58.0% with only a slight increment in parameters. Table 5 showcases the influence of different numbers of shuffling patches. The performance is not proportional to the number of shuffling patches, indicating the importance of moderate content continuity during training.
@@ 5 Conclusion
In this work, we formulate the novel task of Open Panoramic Segmentation (OPS) for unfolding holistic scene understanding, which reaches Open FoV, Open Vocabulary, and Open Domain in surrounding scene segmentation. We have designed the Deformable Adapter Network (DAN) and Random Equirectangular Projection (RERP) for enhancing zero-shot and distortion-aware segmentation capacities of the novel proposed OOOPS model. Extensive experiments on WildPASS, Stanford2D3D, and Matterport3D panoramic scene understanding benchmarks show that the proposed method significantly elevates the performance of state-of-the-art panoramic segmenters.
Limitations. The 360° boundaries of panoramas are not covered in the architectural design, which leaves space for enhancing seamless scene segmentation. Moreover, the generalization capacity of OOOPS could be assessed on surround-view fisheye images. We intend to further extend the proposed solution to panoramic panoptic segmentation.
@@ 0.B.1 Visualization of Segmentation Predictions
0.B.2 Visualization of Deformable Offsets
Figure 11: Visualization of deformable offsets on Stanford2D3D dataset. The green points $\bullet$ are the sample locations. The red points $\bullet$ are deformable offsets in $2$ levels, indicating a deformable receptive field (e.g., each level has a $3\times 3$ kernel size, resulting in $(3\times 3)^{2}=81$ red points). Zoom in for a better view.
Figure 12: Visualization of deformable offsets on Matterport3D dataset. The green points $\bullet$ are the sample locations. The red points $\bullet$ are deformable offsets in $2$ levels, indicating a deformable receptive field (e.g., each level has a $3\times 3$ kernel size, resulting in $(3\times 3)^{2}=81$ red points). Zoom in for a better view.
Appendix 0.C Discussion
@@
Appendix 0.D Acknowledgement
This work was supported in part by the Ministry of Science, Research and the Arts of Baden-Württemberg (MWK) through the Cooperative Graduate School Accessibility through AI-based Assistive Technology (KATE) under Grant BW6-03, and in part by the Helmholtz Association Initiative and Networking Fund on the HAICORE@KIT and HOREKA@KIT partition.

### v2

@@ Abstract
Panoramic images, capturing a 360° field of view (FoV), encompass omnidirectional spatial information crucial for scene understanding. However, it is not only costly to obtain training-sufficient dense-annotated panoramas but also application-restricted when training models in a close-vocabulary setting. To tackle this problem, in this work, we define a new task termed Open Panoramic Segmentation (OPS), where models are trained with FoV-restricted pinhole images in the source domain in an open-vocabulary setting while evaluated with FoV-open panoramic images in the target domain, enabling the zero-shot open panoramic semantic segmentation ability of models. Moreover, we propose a model named OOOPS with a Deformable Adapter Network (DAN), which significantly improves zero-shot panoramic semantic segmentation performance. To further enhance the distortion-aware modeling ability from the pinhole source domain, we propose a novel data augmentation method called Random Equirectangular Projection (RERP) which is specifically designed to address object deformations in advance. Surpassing other state-of-the-art open-vocabulary semantic segmentation approaches, a remarkable performance boost on three panoramic datasets, WildPASS, Stanford2D3D, and Matterport3D, proves the effectiveness of our proposed OOOPS model with RERP on the OPS task, especially +2.2% on outdoor WildPASS and +2.4% mIoU on indoor Stanford2D3D. The source code is publicly available at OPS.
@@ 2.1 Open-vocabulary Semantic Segmentation
Motivated by the success of vision-language pre-training models, Open-Vocabulary Semantic Segmentation (OVSS) has been recently explored to achieve open-set pixel-wise understanding [32, 71, 18, 12, 54, 72, 59, 45, 35, 40, 19, 41, 62, 33, 7, 11, 57, 35, 60, 65, 29, 38]. Most works leverage the open-vocabulary capabilities of CLIP [50] to reason about relations between dense pixel semantics and arbitrary class labels [9, 33, 62, 39, 72, 12]. In [90, 79], multiple datasets are unified by transferring heterogeneous labels into text embeddings for training to achieve OVSS. In [36, 68], image-caption datasets are employed for learning alignment between visual regions and text entities. To reduce the computational costs of using a frozen pre-trained CLIP model for identifying novel classes, latest studies [70, 21] have also proposed efficient architectures for fast OVSS. SAN [70, 69] has designed an asymmetric input resolution strategy to decouple mask proposal prediction and attention bias prediction applied to CLIP for mask class recognition. Furthermore, open-vocabulary panoptic segmentation has been tackled in [12, 67, 8, 13, 81, 46, 72, 67, 35] to render universal image understanding. In this work, we focus on the open-vocabulary zero-shot learning on panoramas instead of pinhole images discussed in the aforementioned works.
@@ 2.3 Deformable Neural Networks
In the realm of deformable neural networks, notable advancements have been made through various iterations. The pioneering work [10] introduces the concept of deformable convolution, laying the foundation for subsequent developments. Building upon this foundation, DCNv2 [91] further refines deformable convolution by introducing a modulation mechanism that expands the scope of deformation modeling. The evolution of deformable neural networks continues with the introduction of DCNv3 [58], which combines the group convolution with DCNv2 for the sake of imitating multi-head mechanism of Transformer [56]. DCNv4 [66] aims to speedup DCNv3, achieving a more efficient model. Additionally, Trans4PASS [83, 84] proposes Deformable Patch Embedding (DPE) and Deformable MLP (DMLP) to handle the distortion that occurs in panoramas. Incorporating the principles of deformable convolution into vision transformers [14], Xia et al. propose DeformableViT [63, 64], extending the applicability of deformable neural networks to the domain of transformer architectures. After rethinking a series of DCN designs, we propose DAO to improve the model’s ability to handle image distortion in the OPS task.
@@ 3.1 Open Panoramic Segmentation
Figure 2: Overview of the OOOPS model architecture. It consists of a frozen CLIP model and a Deformable Adapter Network (DAN) which includes Transformer Layers and the proposed DAO.
@@ 3.3 Deformable Adapter Network
The Deformable Adapter Network is a combination of multiple Transformer [56] Layers and the proposed DAO. Since distortion exists in panoramic images, which is a big challenge when utilizing informative panoramas, we delve deeper into the deformable design [58] and sampling methods like APES [61], MateRobot [86], proposing DAO to tackle the image distortions and object deformations of panoramas. To provide a detailed rethinking process of the deformable convolution design, we introduce the preliminary of the DCN series as follows.
@@
where $\mathbf{m}_{k}$ is the learnable modulation scalar for the $k$-th location. Inspired by Transformer, DCNv3 [58] is proposed with a grouping operation, further boosting the deformation-aware ability of DCNv2. DCNv3 can be addressed with the following formula:
@@
Deformable Adapter Operator (DAO). When dealing with distortions in panoramas, DCNv3 and DCNv4 fall short of meeting the distortion-aware requirements. This is illustrated by the visualization of Fig. 6 in Sec. 4.5. Therefore, DAO is proposed to tackle the distortion problem in panoramic images using the following expression:
@@ 3.4 Random Equirectangular Projection
The Equirectangular Projection [51] (ERP) is one of the most common methods for mapping a globe into a panoramic plane, which transforms spherical coordinates into planar coordinates expressed as follows:
@@ 4.1 Datasets
We train our model using the pinhole general-object COCOStuff-164k [5] while evaluating the model performance on three different panoramic datasets: panoramic street-view WildPASS [77], panoramic indoor-scene Stanford2D3D [2] and panoramic indoor-scene Matterport3D [6].
@@
Matterport3D. The panoramic indoor dataset includes $10,800$ panoramic views from $194,400$ RGB-D images. According to the data preprocessing pipeline of 360BEV [55], a subset of $772$ images in the size of $512{\times}1024$ with $20$ classes is used to evaluate the segmentation performance.
@@ 4.4 Comparison with the State of the Art
Table 1: Results on WildPASS. The mIoU, open mIoU, and per-class IoU are in percentage (%). #PARAMs is the number of learnable parameters in millions.
@@
Table 2: Results on Stanford2D3D. The mIoU, open mIoU, and per-class IoU are in percentage (%). #PARAMs is the number of learnable parameters in millions.
@@
Table 3: Results on Matterport3D. The mIoU, open mIoU, and per-class IoU are in percentage (%). #PARAMs is the number of learnable parameters in millions.
@@ 4.5 Ablation Study
We replace DAO in OOOPS, as illustrated in Fig. 2, with DCN [10], DCNv2 [91], DCNv3 [58] and DCNv4 [66]. When referring to the component substitution without data augmentation in Table 5, it is clear that our DAO surpasses other counterparts with a maximum performance gain of +1.4% mIoU. As for the component substitution with ERP, all models achieve a boost in the range of ${+}0.3\%$ to ${+}0.7\%$ compared with the one without data augmentation. The gain is even bigger considering the component substitution with RERP, achieving ${+}1.0\%$ mIoU. It is worth noting that RERP is able to boost the model performance stably among all models. The best model performance is achieved by the combination of DAO and RERP, yielding 58.0% mIoU with only a slight increment in parameters. Table 5 showcases the influence of different numbers of shuffling patches. The performance is not proportional to the number of patches, indicating the importance of moderate content continuity during training.
@@ 5 Conclusion
In this work, we formulate the novel task of Open Panoramic Segmentation (OPS) for unfolding holistic scene understanding, which reaches Open FoV, Open Vocabulary, and Open Domain in surrounding scene segmentation. We have designed the Deformable Adapter Network (DAN) and Random Equirectangular Projection (RERP) for enhancing distortion-aware segmentation capacities of the novel proposed OOOPS model. Extensive experiments on WildPASS, Stanford2D3D, and Matterport3D benchmarks show that the proposed method significantly elevates the performance of state-of-the-art panoramic segmenters.
Acknowledgements
This work was supported in part by the Ministry of Science, Research and the Arts of Baden-Württemberg (MWK) through the Cooperative Graduate School Accessibility through AI-based Assistive Technology (KATE) under Grant BW6-03, in part by BMBF through a fellowship within the IFI programme of DAAD, in part by the Helmholtz Association Initiative and Networking Fund on the HAICORE@KIT and HOREKA@KIT partition, in part by the National Key RD Program under Grant 2022YFB4701400, and in part by Hangzhou SurImage Technology Company Ltd.
@@ 0.B.2 Visualization of Deformable Offsets
Figure 11: Visualization of deformable offsets on Stanford2D3D dataset.
Figure 12: Visualization of deformable offsets on Matterport3D dataset.
Appendix 0.C Ablation Study
0.C.1 Adaptive Shuffling Patches
We conduct an experiment that divides pinhole images into parts of adaptive sizes when doing RERP augmentation, similar to Mosaic augmentation. The results in Table 7 indicate adaptive sizes can further boost the model performance.
Table 7: mIoU of RERP with adaptive sizes on WildPASS, Stanford2D3D, and Matterport3D datasets. mIoU is in percentage (%).
Method WildPASS Stanford2D3D Matterport3D
RERP 58.0 41.1 31.2
RERP w/ adaptive sizes 58.5 41.5 31.6
0.C.2 Simple Horizontal Rotation
The simple horizontal rotation [31] is used to get a new panorama with a different viewpoint. It is applied to panoramas, not pinhole images. RERP is used to transform a pinhole image into a panorama-like image. It is applied to pinhole images, not panoramas. We experiment with applying simple horizontal rotation to pinhole images. From Table 8 we find that simple horizontal rotation does not bring gains, which falls behind the one with RERP.
Table 8: mIoU of RERP and simple horizontal rotation on WildPASS, Stanford2D3D, and Matterport3D datasets. mIoU is in percentage (%).
Method WildPASS Stanford2D3D Matterport3D
OOOPS w/o RERP 57.0 39.5 31.1
OOOPS w/ Rotation 57.0 39.5 31.1
OOOPS w/ RERP 58.0 41.1 31.2
Appendix 0.D Discussion